Jack is twice as old as Jill.In 2 years,jack will be 4 times as old as Jill was 9 years ago. How old are they know
Answer by checkley71(8405)

let Jack's age=2x & Jill's age=x now.
in 2 years Jack will be (2x+2) years old.
(x-9) was Jill's age 9 years ago.
so we have:
2x+2=4(x-9) or in 2 years Jack will be 4 times as old as Jill was 9 years ago.
2x+2=4x-36
2x-4x=-36-2
-2x=-38
x=-38/-2
x=19 years old is Jill's current age.
seeing as Jack is twice Jill's age then Jack is 2*19=38 years of age.
proof:
2*19+2=4(19-9)
38+2=4*10
40=40